OpenStudy (jaedyn):

When solving negative one over eight (x + 35) = −7, what is the correct sequence of operations?

jhonyy9 (jhonyy9):

1 - ---(x+35) = -7 do you think it in this way ? 8

OpenStudy (brebre123):

are u taking math for college readiness?

jhonyy9 (jhonyy9):

please rewrite it correct

OpenStudy (jaedyn):

Is it multiply each side by negative 1 over 8, add 35 to each each side or multiply each side by negative 1 over 8, add 35 from each side?

OpenStudy (jaedyn):

Anybody?

OpenStudy (shadowlegendx):

\[-\frac{ 1 }{ 8 }(x + 35) = -7\] Does it look like this?

OpenStudy (jaedyn):

Yes

OpenStudy (shadowlegendx):

What do you think you do first?

OpenStudy (jaedyn):

I don't know

OpenStudy (jaedyn):

Can you just help me on this

OpenStudy (shadowlegendx):

Have you ever heard of distribution before?

OpenStudy (jaedyn):

Yes

OpenStudy (jaedyn):

I think I got it

OpenStudy (shadowlegendx):

Mhm, ok

OpenStudy (jaedyn):

Is it multiply each side by -8, subtract 35 from each side?

OpenStudy (shadowlegendx):

Show me your work, you might have this down :)

OpenStudy (jaedyn):